I need help please. How do you form a quadratic equation from this?
If x=-3 and x=4, then form a quadratic equation.

Answer by ankor@dixie-net.com(22740)

How do you form a quadratic equation from this?
If x=-3 and x=4, then form a quadratic equation.
:
If you know how to factor a quadratic equation, then reverse the process:
x = -3 and x = 4, gives us factors of
(x+3)(x-4) = 0
FOIL
x^2 - 4x + 3x - 12 = 0
so we have:
y = x^2 - x - 12